3. Vx+y - 1 Math 2171 hw set 4, page 2 of 2 due October 17
Answer #1

② dy - Suty I - Let U=U12) y=4-8 where dn by ② uno du -) - x +4-1-) du -/ - Ju - du = Tu (separable equation) du - 2 u = x to☺ dy DCreomogeneous e (Homogeneous equalem 4= Vt thiu 312v tal = 1- 48h gos deparable equalra -3 ln 11-4v7 = ent the en 11-48-30 (1-48²) 18 = the (1-442) 316 - te v=ylt Back substitute 1 = + (1 -4 4 2 3 18 G 1 = t [t²_4427/8 let lle =0 Ic- [t² - 44 7